### v2

@@ Abstract
In this paper we propose a novel modification of Contrastive Language-Image Pre-Training (CLIP) guidance for the task of unsupervised backlit image enhancement. Our work builds on the state-of-the-art CLIP-LIT approach, which learns a prompt pair by constraining the text-image similarity between a prompt (negative/positive sample) and a corresponding image (backlit image/well-lit image) in the CLIP embedding space. Learned prompts then guide an image enhancement network. Based on the CLIP-LIT framework, we propose two novel methods for CLIP guidance. First, we show that instead of tuning prompts in the space of text embeddings, it is possible to directly tune their embeddings in the latent space without any loss in quality. This accelerates training and potentially enables the use of additional encoders that do not have a text encoder. Second, we propose a novel approach that does not require any prompt tuning. Instead, based on CLIP embeddings of backlit and well-lit images from training data, we compute the residual vector in the embedding space as a simple difference between the mean embeddings of the well-lit and backlit images. This vector then guides the enhancement network during training, pushing a backlit image towards the space of well-lit images. This approach further dramatically reduces training time, stabilizes training and produces high quality enhanced images without artifacts, both in supervised and unsupervised training regimes. Additionally, we show that residual vectors can be interpreted, revealing biases in training data, and thereby enabling potential bias correction.
@@
Backlit image enhancement Vision-language models
@@ 1 Introduction
Another challenge to high quality automatic backlit image enhancement is the lack of paired data, i.e. backlit images each with a corresponding well-lit image. It is complicated to collect such data, so unsupervised approaches that work with unpaired data are imperative. Recently, CLIP-LIT[13] was proposed, which successfully utilizes CLIP[18] model guidance for training an unsupervised backlit enhancement model. CLIP-LIT uses prompt learning techniques to provide CLIP guidance. More specifically, it constructs two learnable text prompts, which are trained to have CLIP embeddings close to the well-lit and backlit images, respectively. These learnt prompts then guide the image enhancement model during training. Next, several iterations of additional prompt correction followed by enhancement model fine-tuning using the updated prompts are required to produce the final image enhancement model. This approach is the current state-of-the-art in unsupervised backlit image enhancement.
In this work, we demonstrate that prompt training is not the most efficient way to implement the CLIP guidance, and propose two novel methods, named CLIP-LIT-Latent and ResiduAl Vector Embedding (RAVE). In CLIP-LIT-Latent, we train vectors corresponding to well-lit and backlit images directly in the CLIP latent space rather than in text embedding space. We show that this approach has similar or superior performance in both supervised and unsupervised training regimes. Further, in RAVE, we form the guidance vector by subtracting the mean of CLIP latent vectors of backlit images from the mean of CLIP latent vectors of well-lit images. This residual vector points in a direction of moving backlit images to well-lit images in the CLIP embedding space. This vector is then used as guidance for the image enhancement model during training. We show that such guidance enables the enhancement model to produce high quality images, while requiring much less training time. While having all these benefits, both proposed CLIP-LIT-Latent and RAVE retain all the advantages of the original CLIP-LIT method, being lightweight for inference and suitable for both supervised and unsupervised training regimes. Moreover, we show that residual vector used for guidance in RAVE is interpretable, which opens up new possibilities for overcoming biases in the training data.
@@
    We show that these approaches result in similar or better quality according to quantitative metrics in both supervised and unsupervised settings, and superior visual quality of resulting images, while requiring considerably less time to train.
@@ 2 Related work
Backlit Image Enhancement. For the task of backlit image enhancement, many approaches have been proposed. Some techniques aim to identify backlit regions and correct them separately from the rest of the image[12]. Methods based on the Retinex[9] theory divide an image into a reflectance image and an illumination map, and both parts are then enhanced and then combined back together to form the resulting enhanced image. Another approach to enhancing backlit images is using end-to-end deep learning approaches, as in Restormer[25] or Retinexformer[3] methods. Combined techniques based on Retinex theory and deep learning were proposed for both backlit and low-light image enhancement, such as Diff-Retinex[24]. These methods decomposing an input image into a reflectance image and an illumination map and apply diffusion methods for enhancement before combining them to produce the output image. Finally, another set of methods are based on high dynamic range (HDR) imaging which combine multiple exposures to produce a composite well-lit image[19, 27]. However, in this work we only focus on single-image methods where back-lit correction is applied as a post-process.
Fundamentally, backlit image correction is ill-posed, in that for a single backlit image, multiple well-lit variants with high perceptual quality exist. Relatedly, generation of paired training data requires hiring professional photographers, which can be expensive. To help address this problem, recently the BAID[14] dataset was proposed, which contains paired images generated by professional photographers. Nevertheless, development of methods for backlit image enhancement, which can be trained using non-paired data (in an unsupervised manner) remains essential.
@@ 3.1 CLIP-LIT overview
Figure 2: Overview of the original CLIP-LIT approach. (a) depicts the first stage of training, which consists of prompt initialization and the initial training of an enhancement network. (b) shows the second stage, where prompt refinement and enhancement model fine-tuning are iteratively repeated.
(a) Initial Prompts and Enhancement Training
(b) Prompt Refinement and Enhancement Tuning
Figure 3: Overview of the proposed CLIP-LIT-Latent approach. (a) depicts the first stage of training, which consists of latent vectors initialization and the initial training of an enhancement network. (b) shows the second stage, where latent vectors refinement and enhancement model fine-tuning are iteratively repeated.
The CLIP-LIT framework consists of two stages, described below and illustrated on Fig.  2.
@@
$\begin{split}L^{1}_{\text{prompt}}&=\max(0,S(I_{w})-S(I_{b})+m_{0})+max(0,S(I_{t-1})-S(I_{b})+m_{0})\\ &+max(0,S(I_{w})-S(I_{t})+m_{1})+max(0,S(I_{t})-S(I_{t-1})+m_{2})\end{split}$ (5)
@@
Here $m_{0}\in[0,1]$ is the margin between the score of well-lit or enhanced results and the backlit images in the CLIP embedding space; $m_{1}\in[0,1]$ is the margin between the score of the enhanced results and the well-lit images in the CLIP embedding space; $m_{2}\in[0,1]$ is the margin between the newly enhanced results $I_{t}$ and previously enhanced results $I_{t-1}$.
@@ 3.2 CLIP-LIT-Latent
Like CLIP-LIT, CLIP-LIT-Latent also consists of two stages of initial prompts and enhancement training and prompt refinement and enhancement tuning. Fig.  3 illustrates the CLIP-LIT-Latent framework.
@@
Based on these these initialized vectors $E_{p}$ and $E_{n}$, an enhancement network is trained with a CLIP-guided loss. As before, the enhancement model is a Unet which takes backlit image $I_{b}$ as input and outputs an estimated illumination map $I_{i}\in\mathbb{R}^{H\times W\times 1}$. And, as before, it is trained using $L_{\text{enhance}}$ loss, where the $L_{clip}$ component is changed as follows:
@@
CLIP-LIT-Latent has several advantages over CLIP-LIT. First, CLIP-LIT-Latent does not use the text encoder, which speeds up the training and inference processes, as it is no longer necessary to pass vectors and gradients through the text encoder. This also potentially enables the use of other vision models which do not have text encoder parts rather than CLIP for guidance. Second, images obtained from CLIP-LIT-Latent tend to have more contrast and have better visual quality (see Fig. 5 and Fig. 1). We speculate this results from tuning vectors $E_{p}$ and $E_{n}$ directly in the CLIP latent space. In contrast, in the original CLIP-LIT, vectors $T_{p}$ and $T_{n}$ belong to the text embedding space, and must to be tuned so that their projection through the CLIP text encoder satisfies desired conditions. Such optimization is more complex and makes tuning of $T_{p}$ and $T_{n}$ harder and less accurate.
@@ 3.3 RAVE: Residual Vector Embedding
Figure 4: Overview of RAVE model. (a) First, we calculate residual vector $\mathbf{v}_{\text{residual}}$ based on backlit and well-lit training data. (b) Then we switch to enhancement model training based on identity loss and loss based on residual vector.
@@ 4 Experiments
Datasets Following CLIP-LIT, we use the BAID[14] and DIV2K[1] datasets for training. We train our models in two settings: supervised and unsupervised. For the supervised setting, we use 380 backlit and corresponding well-lit images from BAID dataset. For the unsupervised setting, we use 380 backlit images from the BAID and 384 images from DIV2K dataset. For the unsupervised setting, we use the same data as used in CLIP-LIT.
For testing, we follow recent works and use the BAID[14] test dataset, which has 368 backlit and corresponding well-lit images. The data is publicly available.
Training and inference We follow parameter settings of the original CLIP-LIT approach when setting parameters for CLIP-LIT-Latent and RAVE (where applicable), as well as for reproducing CLIP-LIT. Spicific details of training and values of hyperparameters are presented in supplementary material.
@@ 4.1 Results
Quantitative comparison The quantitative comparison on the BAID test dataset is presented in Tab. 1. We report test set metrics of the best model checkpoint obtained during training for all the models. Detailed metrics values for all the checkpoints after each training epoch can be found in supplementary material. Note that many approaches can only be used in supervised setting, that is why their results are only present in the “supervised” section. We see that our method achieves state-of-the-art performance in both settings. In supplementary material, we also present results on time and memory needed to train all the models, and confirm that both CLIP-LIT-Latent and RAVE are much more efficient in that sence.
@@
(a) Visual comparison of results obtained by CLIP-LIT, CLIP-LIT-Latent and RAVE, all trained in supervised setting. RAVE produces results with more general contrast.
(b) Visual comparison of results obtained by CLIP-LIT, CLIP-LIT-Latent and RAVE, all trained in unsupervised setting. RAVE produces results with more general contrast.
Figure 5: Visual comparison of results obtained by CLIP-LIT, CLIP-LIT-Latent and RAVE
@@
Interpretation of the residual vector In RAVE, we build the residual vector $\mathbf{v}_{residual}$ following formula 10. We noted that $\mathbf{v}_{residual}$ is a vector that points to a direction of moving from backlit images to well-lit images in the CLIP embedding space. Here we prove this by interpreting vector $\mathbf{v}_{residual}$ using the fact that CLIP has a common vector space for embeddings of images and text.
We go through the vocabulary of CLIP text encoder and find those vocabulary tokens, which have the closest and farthest CLIP embeddings to the $\mathbf{v}_{residual}$ in the CLIP latent space in terms of cosine similarity. More formally, we use the following metric:
@@
Here $e_{token}$ is embedding of vocabulary $token$, which is input to the CLIP text encoder $\Phi_{text}$. We sort vocabulary tokens using $sim$. Table 2 shows 7 vocabulary tokens which have highest and lowest cosine similarity to the $v_{residual}$ vectors obtained from paired dataset used for supervised training and from unpaired dataset used in unsupervised setting.
@@
For the supervised setting, many of the least similar tokens have a meaning around "darkness": "silhouette", "dark", "shadow", etc., and absolute value of their cosine similarity to $v_{\text{residual}}$ is around $0.12-0.15$. At the same time, most similar tokens do not share any meaningful semantics, and the absolute value of their cosine similarity to $v_{\text{residual}}$ is much smaller, being around $0.01-0.04$. This supports the hypothesis that $v_{\text{residual}}$ has a meaning of shifting from "dark", "shadowy" images to "normal" images, having average level of lightness. $v_{\text{residual}}$ does not carry any other semantic meaning, as it was calculated using paired dataset , where for each backlit image there is its well-lit analogue, and all the semantics present in the BAID dataset is cancelled as a result of subtraction of the mean well-lit and backlit vectors when constructing $v_{\text{residual}}$. On the other hand, for the unsupervised setting, we see that the least similar tokens have a meaning around "asian": "busan", "beijing", etc., and the most similar ones have a meaning around "nature": "wildlife", "southafrica", "countrylife", etc. Herein, an absolute value of cosine similarities for both sets of tokens if quite high, being around $0.08-0.12$. This reflects the bias in the datasets used in unsupervised setting. Indeed, in the BAID dataset there are mostly images of Chinese cities or people, while in DIV2K dataset there are many pictures of nature. Due to the bias, RAVE training in the unsupervised setting becomes less effective, making the enhancement model shift images not in an ideal direction in the CLIP latent space. Nevertheless, RAVE still shows good performance in unsupervised setting. Based on this analysis, $v_{\text{residual}}$ is effective in finding biases in the datasets and creating techniques of mitigating them.
@@
Table 2: Vocabulary tokens and their cosine similarities, which have the lowest and highest cosine similarity to the $v_{\text{residual}}$ vectors calculated for supervised and unsupervised training datasets.
Supervised Unsupervised
@@
Necessity of the residual vector Given the interpretation of $\mathbf{v}_{\text{residual}}$, we investigate whether it is necessary to compute $\mathbf{v}_{\text{residual}}$ using the training data, or is it enough to use CLIP latent vectors of words such as "dark", "darkness" as a guidance vector for enhancement model. We trained RAVE using $v_{\text{residual}}=\Phi_{\text{text}}(token)$ as guidance vector, where $token\in\{"silhouette","dark","darkness"\}$, but did not observe any meaningful enhancement in the resulting images (examples of such resulting images can be found in supplementary material). We suspect that this is due to the fact that latent vectors of backlit and well-lit images have similar cosine similarity to embeddings of these tokens, as presented in Tab 3. This suggests that residual vector $\mathbf{v}_{\text{residual}}$ is not simply a combination of latent vectors of some tokens, but it contains more specific information with the meaning of shifting from backlit to well-lit data.

### v3

@@ Abstract
In this paper we propose a novel modification of Contrastive Language-Image Pre-Training (CLIP) guidance for the task of backlit image enhancement. Our work builds on the state-of-the-art CLIP-LIT approach, which learns a prompt pair by constraining the text-image similarity between a prompt (negative/positive sample) and a corresponding image (backlit image/well-lit image) in the CLIP embedding space. Learned prompts then guide an image enhancement network. Based on the CLIP-LIT framework, we propose two novel methods for CLIP guidance. First, we show that instead of tuning prompts in the space of text embeddings, it is possible to directly tune their embeddings in the latent space without any loss in quality. This accelerates training and potentially enables the use of additional encoders that do not have a text encoder. Second, we propose a novel approach that does not require any prompt tuning. Instead, based on CLIP embeddings of backlit and well-lit images from training data, we compute the residual vector in the embedding space as a simple difference between the mean embeddings of the well-lit and backlit images. This vector then guides the enhancement network during training, pushing a backlit image towards the space of well-lit images. This approach further dramatically reduces training time, stabilizes training and produces high quality enhanced images without artifacts. Additionally, we show that residual vectors can be interpreted, revealing biases in training data, and thereby enabling potential bias correction. Code is available at https://github.com/Atmyre/RAVE
@@
Backlit image enhancement Vision-language models Residual vector embedding
Accepted to ECCV 2024
@@ 1 Introduction
Another challenge to high quality automatic backlit image enhancement is the lack of paired data, i.e. backlit images each with a corresponding well-lit image. It is complicated to collect such data, so approaches that work with unpaired data are imperative. Recently, CLIP-LIT[14] was proposed, which successfully utilizes CLIP[19] model guidance for training a backlit enhancement model with unpaired data. CLIP-LIT uses prompt learning techniques to provide CLIP guidance. More specifically, it constructs two learnable text prompts, which are trained to have CLIP embeddings close to the well-lit and backlit images, respectively. These learnt prompts then guide the image enhancement model during training. Next, several iterations of additional prompt correction followed by enhancement model fine-tuning using the updated prompts are required to produce the final image enhancement model. This approach is the current state-of-the-art in backlit image enhancement with training on unpaired data.
In this work, we demonstrate that prompt training is not the most efficient way to implement the CLIP guidance, and propose two novel methods, named CLIP-LIT-Latent and ResiduAl Vector Embedding (RAVE). In CLIP-LIT-Latent, we train vectors corresponding to well-lit and backlit images directly in the CLIP latent space rather than in text embedding space. We show that this approach has similar or superior performance when trained on paired or unpaired image data. Further, in RAVE, we form the guidance vector by subtracting the mean of CLIP latent vectors of backlit images from the mean of CLIP latent vectors of well-lit images. This residual vector points in a direction of moving backlit images to well-lit images in the CLIP embedding space. This vector is then used as guidance for the image enhancement model during training. We show that such guidance enables the enhancement model to produce high quality images, while requiring much less training time. While having all these benefits, both proposed CLIP-LIT-Latent and RAVE retain all the advantages of the original CLIP-LIT method, being lightweight for inference and suitable for training on paired or unpaired data. Moreover, we show that residual vector used for guidance in RAVE is interpretable, which opens up new possibilities for overcoming biases in the training data.
@@
    We show that these approaches result in similar or better quality according to quantitative metrics in both settings of training on paired or unpaired data, and superior visual quality of resulting images, while requiring considerably less time to train.
@@ 2 Related work
Backlit Image Enhancement. For the task of backlit image enhancement, many approaches have been proposed. Some techniques aim to identify backlit regions and correct them separately from the rest of the image[13]. Methods based on the Retinex[10] theory divide an image into a reflectance image and an illumination map, and both parts are then enhanced and then combined back together to form the resulting enhanced image. Another approach to enhancing backlit images is using end-to-end deep learning approaches, as in Restormer[26] or Retinexformer[3] methods. Combined techniques based on Retinex theory and deep learning were proposed for both backlit and low-light image enhancement, such as Diff-Retinex[25]. These methods decomposing an input image into a reflectance image and an illumination map and apply diffusion methods for enhancement before combining them to produce the output image. Finally, another set of methods are based on high dynamic range (HDR) imaging which combine multiple exposures to produce a composite well-lit image[20, 28, 4]. However, in this work we only focus on single-image methods where back-lit correction is applied as a post-process.
Fundamentally, backlit image correction is ill-posed, in that for a single backlit image, multiple well-lit variants with high perceptual quality exist. Relatedly, generation of paired training data requires hiring professional photographers, which can be expensive. To help address this problem, recently the BAID[15] dataset was proposed, which contains paired images generated by professional photographers. Nevertheless, development of methods for backlit image enhancement, which can be trained using non-paired data remains essential.
@@ 3.1 CLIP-LIT overview
Figure 2: Overview of the original CLIP-LIT approach and proposed CLIP-LIT-Latent. (a) depicts the first stage of training, which consists of prompt or latent vector initialization and the initial training of an enhancement network. (b) shows the second stage, where prompt/latent vector refinement and enhancement model fine-tuning are iteratively repeated. Blue and red boxes are related to CLIP-LIT and CLIP-LIT-Latent.
The CLIP-LIT framework (Fig.  2) consists of two stages, described below.
@@
$\begin{split}L^{1}_{\text{prompt}}&=\max(0,S(I_{w})-S(I_{b})+m_{0})+\max(0,S(I_{t-1})-S(I_{b})+m_{0})\\ &+\max(0,S(I_{w})-S(I_{t})+m_{1})+\max(0,S(I_{t})-S(I_{t-1})+m_{2})\end{split}$ (5)
@@
Here $m_{0}\in[0,1]$ is the margin between the score of well-lit or enhanced result and the backlit image in the CLIP embedding space; $m_{1}\in[0,1]$ is the margin between the score of the enhanced result and the well-lit image in the CLIP embedding space; $m_{2}\in[0,1]$ is the margin between the newly enhanced result $I_{t}$ and previously enhanced result $I_{t-1}$.
@@ 3.2 CLIP-LIT-Latent
Like CLIP-LIT, CLIP-LIT-Latent also consists of two stages of initial prompts and enhancement training and prompt refinement and enhancement tuning. Fig.  2 illustrates the CLIP-LIT-Latent framework.
@@
Based on these initialized vectors $E_{p}$ and $E_{n}$, an enhancement network is trained with a CLIP-guided loss. As before, the enhancement model is a Unet which takes backlit image $I_{b}$ as input and outputs an estimated illumination map $I_{i}\in\mathbb{R}^{H\times W\times 1}$. And, as before, it is trained using $L_{\text{enhance}}$ loss, where the $L_{\text{clip}}$ component is changed as follows:
@@
CLIP-LIT-Latent has several advantages over CLIP-LIT. First, CLIP-LIT-Latent does not use the text encoder, which speeds up the training and inference processes, as it is no longer necessary to pass vectors and gradients through the text encoder. This also potentially enables the use of other vision models which do not have text encoder parts rather than CLIP for guidance. Second, images obtained from CLIP-LIT-Latent tend to have more contrast and have better visual quality (see Fig. 4 and Fig. 1). We speculate this results from tuning vectors $E_{p}$ and $E_{n}$ directly in the CLIP latent space. In contrast, in the original CLIP-LIT, vectors $T_{p}$ and $T_{n}$ belong to the text embedding space, and must be tuned so that their projection through the CLIP text encoder satisfies desired conditions. Such optimization is more complex and makes tuning of $T_{p}$ and $T_{n}$ harder and less accurate.
@@ 3.3 RAVE: Residual Vector Embedding
Figure 3: Overview of the RAVE model. (a) First, we calculate residual vector $\mathbf{v}_{\text{residual}}$ based on backlit and well-lit training data. (b) Then we switch to enhancement model training based on the identity loss and the loss based on the residual vector.
@@ 4 Experiments
Datasets Following CLIP-LIT, we use the BAID[15] and DIV2K[1] datasets for training. We train our models in two settings: using paired and unpaired training data. For the paired setting, we use 380 backlit and corresponding well-lit images from BAID dataset. For the unpaied setting, we use 380 backlit images from the BAID and 384 images from DIV2K dataset. For the unpaired setting, we use the same data as used in CLIP-LIT. For testing, we follow recent literature and use the BAID[15] test dataset, which has 368 backlit and corresponding well-lit images. The data is publicly available.
Training and inference We follow parameter settings of the original CLIP-LIT approach when setting parameters for CLIP-LIT-Latent and RAVE (where applicable), as well as for reproducing CLIP-LIT. Specific details of training and values of hyperparameters are presented in supplementary material.
@@ 4.1 Results
Quantitative comparison The quantitative comparison on the BAID test dataset is presented in Tab. 1. We report test set metrics of the best model checkpoint obtained during training for all the models. Detailed metrics values for all the checkpoints after each training epoch can be found in supplementary material. Note that many approaches can only be trained using paired data, that is why their results are only present in the “paired” section. We see that our method achieves state-of-the-art performance in both settings. In supplementary material, we also present results on time needed to train all the models, and confirm that both CLIP-LIT-Latent and RAVE are much more efficient. Note that in the setting of unpaired data we have two versions of RAVE, referred as RAVE and RAVE shifted in Tab. 1. RAVE shifted is a variant of RAVE with a simple modification of residual vector, see “Interpretation of the residual vector” section below for more details.
@@
RAVE shifted (ours) 21.37 0.877 0.149 40.90
(a) Visual comparison of results by CLIP-LIT, CLIP-LIT-Latent and RAVE, all trained with paired data. RAVE produces results with more general contrast.
(b) Visual comparison of results obtained by CLIP-LIT, CLIP-LIT-Latent and RAVE (shifted version), all trained with unpaired data. RAVE produces results with more general contrast.
Figure 4: Visual comparison of results by CLIP-LIT, CLIP-LIT-Latent and RAVE.
@@
Interpretation of the residual vector In RAVE, we build the residual vector $\mathbf{v}_{residual}$ following formula 16. We noted that $\mathbf{v}_{residual}$ is a vector that points to a direction of moving from backlit images to well-lit images in the CLIP embedding space. Here we prove this by interpreting vector $\mathbf{v}_{residual}$ using the fact that CLIP has a common vector space for embeddings of images and text.
We go through the vocabulary of the CLIP text encoder and find those vocabulary tokens which have the closest and farthest CLIP embeddings to the $\mathbf{v}_{residual}$ in the CLIP latent space in terms of cosine similarity. More formally, we use the following metric:
@@
Here $e_{token}$ is embedding of vocabulary $token$, which is input to the CLIP text encoder $\Phi_{text}$. We sort vocabulary tokens using $sim$. Table 2 shows 7 vocabulary tokens which have highest and lowest cosine similarity to the $v_{residual}$ vectors obtained from the paired dataset and from the unpaired dataset.
@@
In the case of paired data, many of the least similar tokens have a meaning around "darkness": "silhouette", "dark", "shadow", etc., and absolute value of their cosine similarity to $v_{\text{residual}}$ is around $0.12-0.15$. At the same time, most similar tokens do not share any meaningful semantics, and the absolute value of their cosine similarity to $v_{\text{residual}}$ is much smaller, being around $0.01-0.04$. This supports the hypothesis that $v_{\text{residual}}$ has a meaning of shifting from "dark", "shadowy" images to "normal" images, having average level of lightness. $v_{\text{residual}}$ does not carry any other semantic meaning, as it was calculated using paired dataset, where for each backlit image there is its well-lit analogue, and all the semantics present in the BAID dataset is cancelled as a result of subtraction of the mean well-lit and backlit vectors when constructing $v_{\text{residual}}$. On the other hand, in the case of unpaired data, we see that the least similar tokens have a meaning around "asian": "busan", "beijing", etc., and the most similar ones have a meaning around "nature": "wildlife", "southafrica", "countrylife", etc. Herein, an absolute value of cosine similarities for both sets of tokens if quite high, being around $0.08-0.12$. This reflects the bias in the unpaired datasets. Indeed, in the BAID dataset there are mostly images of Chinese cities or people, while in DIV2K dataset there are many pictures of nature. Due to the bias, RAVE training with unpaired data becomes less effective, making the enhancement model shift images not in an ideal direction in the CLIP latent space. Nevertheless, RAVE still shows good performance when trained with unpaired data.
We also propose a simple modification for correcting the residual vector in the unpaired setting. We simply take $n$ most and least similar tokens to form an additional residual vector as follows:
$\begin{split}&\mathbf{v}_{\text{add\_residual}}=f_{\text{norm}}(\mathbf{v}_{\text{closest}}-\mathbf{v}_{\text{farthest}})\\ &\mathbf{v}_{\text{closest}}=f_{norm}\left(\frac{\sum_{i=1}^{n}f_{\text{norm}}(\Phi_{\text{text}}(T^{c}_{i}))}{n}\right)\\ &\mathbf{v}_{\text{farthest}}=f_{norm}\left(\frac{\sum_{i=1}^{n}f_{\text{norm}}(\Phi_{\text{text}}(T^{f}_{i}))}{n}\right)\end{split}$ (15)
$T^{c}_{i}$ and $T^{f}_{i}$ are tokens with $i^{th}$ most and least similar embeddings to the $\mathbf{v}_{\text{residual}}$. And finally we update the $\mathbf{v}_{\text{residual}}$ as follows:
$\mathbf{v}_{\text{residual}}=\mathbf{v}_{\text{residual}}-cos(\mathbf{v}_{\text{residual}},\mathbf{v}_{\text{add\_residual}})\cdot\mathbf{v}_{\text{add\_residual}}$ (16)
Resulting vector $\mathbf{v}_{\text{residual}}$ will keep the meaning of shifting from "dark", images to "normal" images while losing the unwanted semantics. Experimental results show that RAVE trained with such updated $\mathbf{v}_{\text{residual}}$ produces better results (see RAVE shifted in Tab. 1). In supplementary material we present visual comparison of results obtained by original and shifted versions of RAVE, where it can be clearly seen that RAVE shifted produces more well-lit results, while keeping all the advantages of RAVE. There we also present analysis of tokens with the most and least similar embeddings to the updated $\mathbf{v}_{\text{residual}}$. This analysis supports the claim about the meaning of updated $\mathbf{v}_{\text{residual}}$.
@@
Table 2: Vocabulary tokens and their cosine similarities, which have the lowest and highest cosine similarity to the $v_{\text{residual}}$ vectors calculated for paired and unpaired training datasets.
Paired Unpaired
@@
[…]